I'm confused. can you help me find the outputs for f(x)=3x^2 -4x?
f(0) =?
f (-1) =?
f (2) = ?
I understand that what inpute= goes into the () but, it's confusing.

Answer by sudhanshu_kmr(1152)


f(0)= 0
f(-1) = 3 - 4(-1) = 3 + 4 = 7
f(2) = 12 - 8 = 4